OpenStudy (anonymous):

URGENT! PLEASE SOMEBODY HELP!! Factorize and solve the X: x(x-2)=5(x+12)

OpenStudy (anonymous):

first you have to unfactor (aka multiply) to get \[x^2-2x=5x+60\]

OpenStudy (anonymous):

then \[x^2-7x-60=0\] \[(x+5)(x-12)=0\] \[x=-5,x=12\]

OpenStudy (anonymous):

THANK YOU! would you help me with other three maths too? please?